# ATOM: A Pretrained Neural Operator for Multitask Molecular Dynamics

### v1

@@ Abstract
Molecular dynamics (MD) simulations underpin modern computational drug discovery, materials science, and biochemistry. Recent machine learning models provide high-fidelity MD predictions without the need to repeatedly solve quantum mechanical forces, enabling significant speedups over conventional pipelines. Yet many such methods typically enforce strict equivariance and rely on sequential rollouts, thus limiting their flexibility and simulation efficiency. They are also commonly single-task, trained on individual molecules and fixed timeframes, which restricts generalization to unseen compounds and extended timesteps. To address these issues, we propose Atomistic Transformer Operator for Molecules (ATOM), a pretrained transformer neural operator for multitask molecular dynamics. ATOM adopts a quasi-equivariant design that requires no explicit molecular graph and employs a temporal attention mechanism, allowing for the accurate parallel decoding of multiple future states. To support operator pretraining across chemicals and timescales, we curate TG80, a large, diverse, and numerically stable MD dataset with over 2.5 million femtoseconds of trajectories across 80 compounds. ATOM achieves state-of-the-art performance on established single-task benchmarks, such as MD17, RMD17 and MD22. After multitask pretraining on TG80, ATOM shows exceptional zero-shot generalization to unseen molecules across varying time horizons. We believe ATOM represents a significant step toward accurate, efficient, and transferable molecular dynamics models.
@@ 1 Introduction
Molecular dynamics (MD) serves as a computational microscope of atomic motion and is now integral to drug discovery and materials science pipelines (Dror et al., 2012; De Vivo et al., 2016). In ab initio molecular dynamics, quantum-mechanical density functional theory (DFT) is used to compute atomic forces, and the resulting equations of motion are integrated to generate high-fidelity trajectories. However, DFT’s computational complexity scales at least cubically with the number of atoms, and relies on double-precision arithmetic that limits GPU acceleration (Kresse & Furthmüller, 1996; Stein et al., 2020; Li et al., 2024).
Neural approaches have recently emerged as a promising solution to the scalability bottleneck. Equivariant architectures, in particular, encode physical symmetries to model interatomic dynamics, achieving ab initio-level accuracy at significantly reduced computational cost (Batzner et al., 2022; Musaelian et al., 2022; Batatia et al., 2022, 2023; Xu et al., 2024). While equivariance is often deemed essential for improving generalization, strict symmetry preservation involves substantial tradeoffs (Xu et al., 2024; Schreiner et al., 2023). Architectures that enforce exact equivariance at every layer often increase computational overhead, restrict model expressivity, and complicate optimization (Fuchs et al., 2020; Brehmer et al., 2023; Elhag et al., 2025). It is unclear whether symmetry constraints can be relaxed without sacrificing accuracy for molecular dynamics.
Furthermore, most existing methods for molecular dynamics are autoregressive, predicting the next state based on the current one (Köhler et al., 2019; Fuchs et al., 2020; Thiemann et al., 2025). Autoregressive approaches often struggle to capture long-horizon temporal dependencies and accumulate error as the prediction horizon grows (Bengio et al., 2015; Bergsma et al., 2023; Taieb & Atiya, 2016). Inference speeds are also constrained by the need for sequential integration, failing to exploit modern, highly parallel compute architectures. One exception is Equivariant Graph Neural Operator (EGNO) (Xu et al., 2024), which models the entire trajectory with neural operator learning. Nevertheless, EGNO enforces strict equivariance and is single-task in nature, i.e., it is trained and evaluated on trajectories of each molecule separately with a fixed time horizon, which limits zero-shot generalization to unseen molecules or timeframes.
@@
We believe our work represents a shift in molecular dynamics modeling, where we demonstrate the potential of quasi-equivariance designs and zero-shot generalization to out-of-domain systems, which is enabled by the comprehensive TG80 MD dataset.
@@ 2 Related work
Neural Operators. Neural operators are deep learning methods for learning operators between function spaces (Kovachki et al., 2021). A wide variety of architectures have been proposed for such operator learning. Notably, Fourier Neural Operator (FNO) (Li et al., 2021) learns an operator in the Fourier domain, while its derivatives G-FNO (Helwig et al., 2023) and PINO (Li et al., 2023b), respectively, add group equivariance and physics-informed properties. Xu et al. (2024) bridges this framework with molecular dynamics by recasting the task as learning a propagation operator that evolves historical atomic positions into their future configurations. Specifically, EGNO (Xu et al., 2024) is proposed by integrating EGNN and FNO layers to learn dynamic trajectories, capturing both spatial and temporal correlations. Recently, transformer neural operators (Bryutkin et al., 2024; Hao et al., 2023; Li et al., 2023a) have surpassed the performance of FNO in most partial differential equation (PDE) tasks. Notably, OFormer (Li et al., 2023a) uses a linear Galerkin-type attention mechanism, which omits the softmax and instead interprets the latent column vectors as basis functions. General Neural Operator Transformer (GNOT) (Hao et al., 2023) employs a novel subquadratic cross-attention methodology to integrate multiple feature types (e.g., shape and point relationships) into their transformer blocks. With ATOM, we unify the MD problem formulation and temporal discretization approach introduced by EGNO with the increased representational power of transformers in operator settings.
@@
MD Benchmarks. Research on graph machine learning for molecular dynamics suffers from poor benchmarking (Bechler-Speicher et al., 2025). For example, despite the fact that MD17 Benzene exhibits non-physical noise approximately 1000 times higher compared to other compounds (Christensen & von Lilienfeld, 2020), it is still regularly employed to benchmark new models (Bihani et al., 2023; Huang et al., 2022; Liao & Smidt, 2023; Xu et al., 2024). The practical relevance of single-task learning on these datasets is also dubious, as predicting trajectories for molecules with existing numerical solutions offers minimal benefit. We believe the strengths of neural approaches emerge in transfer learning, where models generalize to unseen compounds, thereby circumventing the computational costs associated with explicit numerical simulations. This motivates our development of TG80 to facilitate multitask dynamics learning across molecular systems.
@@ 3.2 ATOM Model and Training Design
Here we outline the pipeline of ATOM. At its core is an equivariant lifting layer (Section 3.2.1), which maps atomic positions, velocities and their phase features into a richer embedding space while preserving symmetry under the Euclidean group $E(3)$. The lifted embeddings are then processed by the ATOM attention block, which applies heterogeneous attention over positions, velocities, and phase features with chemical augmentation (Section 3.2.2). To capture temporal dynamics, we incorporate a temporal rotary position embedding (T-RoPE) (Section 3.2.2) that depends only on time lags and is shared across atoms, ensuring translation invariance in time and permutation invariance within each molecule. Finally, to counter numerical noise in training trajectories, we inject randomly sampled position and velocity perturbations during training (Section 3.2.3), which improves robustness and acts as a regularizer against overfitting. The overall pipeline of ATOM is in Figure 1.
@@ 3.2.1 E(3) Equivariant Lifting
To model atomic states in a symmetry-respecting way, each atom is encoded with its 3D position and velocity, augmented with their norms: $\mathbf{x}=(x,y,z,\sqrt{x^{2}+y^{2}+z^{2}}),\,\mathbf{v}=(v_{x},v_{y},v_{z},\sqrt{v_{x}^{2}+v_{y}^{2}+v_{z}^{2}})$. To construct higher-dimensional features that remain consistent with $E(3)$ symmetry, we apply equivariant lifting that maps the inputs through learnable functions that preserve group actions. Specifically, we use $E(3)$-equivariant linear layers (Geiger & Smidt, 2022) that lifts the position and velocity vectors to a feature space. The resulting features satisfy the equivariance constraints by construction. We further construct phase space of each atom by augmenting the position and velocity vectors with atomic number, which is subsequently processed by a learnable equivariant layer to obtain a lifted representation. The final lifted embedding for a molecule is given by $({\mathbf{X}},{\mathbf{V}},{\mathbf{Z}})\in\mathbb{R}^{3\times NP\times d_{v}}$ corresponding to position, velocity and phase features. The second dimension aggregates nodes and time for attention and $d_{v}$ is the embedding space dimension.
We highlight that after the equivariant lifting layer, we do not enforce equivariance in the subsequent Transformer blocks. This relaxation leads to improved performance compared to fully equivariant designs and, show robustness to random rotations of the trajectories compared to non-equivariant baselines (see Section 4.4).
@@ 3.2.2 ATOM Heterogeneous Temporal Attention
We employ a heterogeneous temporal attention mechanism to enable mixing between multiple features $({\mathbf{X}},{\mathbf{V}},{\mathbf{Z}})\in\mathbb{R}^{3\times NP\times d_{v}}$ across spatial and temporal dimensions. We use the phase space embedding $\mathbf{Z}$} as the query and attend to the key-value pairs formed from all features ${\mathbf{X}},{\mathbf{V}},{\mathbf{Z}}\in\mathbb{R}^{NP\times d_{v}}$. In Figure 7, we show that this improves performance by 6.36% over standard self-attention for single-task prediction. In addition, to encode temporal information, we introduce Temporal RoPE (T-RoPE), an adaptation of RoPE (Su et al., 2023) to irregular time lags by driving the phases with timestamps built from per-step increments $\{\Delta t\}$.
Let the hidden dimension per head be $d_{h}$ (even). We define frequencies $\omega_{k}=b^{-2k/d_{h}}$ for $k=0,\dots,d_{h}/2-1$. Given per-step time increments $\{\Delta t_{p}\}_{p=1}^{P}$, we build timestamps $t_{p}=t+\sum_{r=1}^{p}\Delta t_{r}$, and assign a single rotation to all $N$ atoms at timestep $p$: $\mathbf{R}_{p}=\operatorname{diag}\left(\mathbf{R}(\theta_{p,0}),\dots,\mathbf{R}(\theta_{p,d_{h}/2-1})\right)\in\mathbb{R}^{d_{h}\times d_{h}}$, where $\theta_{p,k}=\frac{\omega_{k}}{\tau}\,\bigl(t_{p}-t_{0}\bigr)$ and $\mathbf{R}(\theta)\in\mathbb{R}^{2\times 2}$ is the rotation matrix with angle $\theta$ and $\tau>0$ is a timescale hyperparameter. Suppose the query molecule state at time $p$ is given as $\mathbf{Q}_{p}\in\mathbb{R}^{N\times d_{h}}$ and key molecule state at time $p^{\prime}$ is $\mathbf{K}_{p^{\prime}}\in\mathbb{R}^{N\times d_{h}}$. We apply $\mathbf{R}_{p},\mathbf{R}_{p^{\prime}}$ to $\mathbf{Q}_{p},\mathbf{K}_{p^{\prime}}$ respectively so that the rotary dot product $\mathbf{Q}_{p}\mathbf{R}_{p}(\mathbf{K}_{p^{\prime}}\mathbf{R}_{p^{\prime}})^{\top}$ depends only on the time interval $t_{p^{\prime}}-t_{p}$. This makes attention translation invariant in time, which allows for interpolation and extrapolation across irregular increments $\{\Delta t_{p}\}$. In addition, sharing the same $\mathbf{R}_{p}$ across all $N$ atoms in a molecule ensures permutation-invariance within a timestep. For aggregated query and key matrices $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{NP\times d_{h}}$, we denote the application of temporal Rotary Position Embedding (RoPE) across $P$ timesteps and $N$ atoms as $\operatorname{T-RoPE}(\mathbf{Q}),\operatorname{T-RoPE}(\mathbf{K})\in\mathbb{R}^{NP\times d_{h}}$.
@@
where $Q(\cdot),K(\cdot),V(\cdot)$ represent the query, key and value projections. We introduce learnable weights $\gamma_{\mathbf{F}}$ to modulate the relative importance of each feature. In Section G.1, we show that heterogeneous attention is equivalent to a kernel integral operator and discuss the properties of the kernel.

### v2

@@ Abstract
Molecular dynamics (MD) simulations underpin modern computational drug discovery, materials science, and biochemistry. Recent machine learning models provide high-fidelity MD predictions without the need to repeatedly solve quantum mechanical forces, enabling significant speedups over conventional pipelines. Yet many such methods typically enforce strict equivariance and rely on sequential rollouts, thus limiting their flexibility and simulation efficiency. They are also commonly single-task, trained on individual molecules and fixed timeframes, which restricts generalization to unseen compounds and extended timesteps. To address these issues, we propose Atomistic Transformer Operator for Molecules (ATOM), a pretrained transformer neural operator for multitask molecular dynamics. ATOM adopts a quasi-equivariant design that requires no explicit molecular graph and employs a temporal attention mechanism, enabling accurate parallel decoding of multiple future states. To support operator pretraining across chemicals and timescales, we curate TG80, a large, diverse, and numerically stable MD dataset with over 2.5 million femtoseconds of trajectories across 80 compounds. ATOM achieves state-of-the-art performance on established single-task benchmarks, such as MD17, RMD17 and MD22. After multitask pretraining on TG80, ATOM shows exceptional zero-shot generalization to unseen molecules across varying time horizons. We believe ATOM represents a significant step toward accurate, efficient, and transferable molecular dynamics models.
@@ 1 Introduction
Molecular dynamics (MD) serves as a computational microscope of atomic motion and is now integral to drug discovery and materials science pipelines (Dror et al., 2012; De Vivo et al., 2016). In ab initio molecular dynamics, quantum-mechanical density functional theory (DFT) is used to compute atomic forces, and the resulting equations of motion are integrated to generate high-fidelity trajectories. However, DFT’s computational complexity scales at least cubically with the number of atoms, and relies on double-precision arithmetic that limits GPU acceleration (Kresse and Furthmüller, 1996; Stein et al., 2020; Li et al., 2024).
Neural approaches have recently emerged as a promising solution to the scalability bottleneck. Equivariant architectures, in particular, encode physical symmetries to model interatomic dynamics, achieving ab initio-level accuracy at significantly reduced computational cost (Batzner et al., 2022; Musaelian et al., 2022; Batatia et al., 2022; 2023; Xu et al., 2024). While equivariance is often deemed essential for improving generalization, strict symmetry preservation involves substantial tradeoffs (Xu et al., 2024; Schreiner et al., 2023). Architectures that enforce exact equivariance at every layer often increase computational overhead, restrict model expressivity, and complicate optimization (Fuchs et al., 2020; Brehmer et al., 2023; Elhag et al., 2025). It is unclear whether symmetry constraints can be relaxed without sacrificing accuracy for molecular dynamics.
Furthermore, most existing methods for molecular dynamics are autoregressive, predicting the next state based on the current one (Köhler et al., 2019; Fuchs et al., 2020; Thiemann et al., 2025). Autoregressive approaches often struggle to capture long-horizon temporal dependencies and accumulate error as the prediction horizon grows (Bengio et al., 2015; Bergsma et al., 2023; Taieb and Atiya, 2016). Inference speeds are also constrained by the need for sequential integration, failing to exploit modern, highly parallel compute architectures. One exception is Equivariant Graph Neural Operator (EGNO) (Xu et al., 2024), which models the entire trajectory with neural operator learning. Nevertheless, EGNO enforces strict equivariance and is single-task in nature, i.e., it is trained and evaluated on trajectories of each molecule separately with a fixed time horizon, which limits zero-shot generalization to unseen molecules or timeframes.
@@
We believe our work represents a shift in molecular dynamics modeling, where we demonstrate the potential of quasi-equivariance designs and zero-shot generalization to out-of-domain systems, which is enabled by the comprehensive TG80 MD dataset. ATOM and TG80 are available at this repository.
@@ 2 Related work
Time-coarsened Molecular Dynamics Time coarsening is a coarse-graining method which preserves molecular structure, but compresses many short integration steps into a few large-stride updates to reduce the cost of long-time simulation (Kmiecik et al., 2016). Stochastic coarse-graining approaches often learn transition kernels on configuration space, bypassing explicit integration of the equations of motion. Klein et al. (2023) learns such a kernel with a normalizing flow and uses it as an MCMC proposal targeting the Boltzmann distribution, Hsu et al. (2024) uses a conditional diffusion model to learn a transition probability matrix, and Yu et al. (2025) uses flow-matching to learn a vector field transporting current states to future states. Closer to our framework, deterministic methods such as MDNet (Zheng et al., 2021) and TrajCast (Thiemann et al., 2025) learn a GNN and EGNN, respectively, which autoregressively predict fixed strides 10-100 times larger than those of MD integrators. Bigi et al. (2025) incorporates Hamiltonian structure and explicit energy-conservation. Most of the methods require direct force learning and are sequential in nature, while ATOM may be interpreted as a force-free deterministic coarse-graining approach, wherein temporal pushforward is approximated by a learned propagation operator which is decoded in parallel.
@@
Neural Operators. Neural operators are deep learning methods for learning operators between function spaces (Kovachki et al., 2021). A wide variety of architectures have been proposed for such operator learning. Notably, Fourier Neural Operator (FNO) (Li et al., 2021) learns an operator in the Fourier domain, while its derivatives G-FNO (Helwig et al., 2023) and PINO (Li et al., 2023b), respectively, add group equivariance and physics-informed properties. Xu et al. (2024) bridges this framework with molecular dynamics by recasting the task as learning a propagation operator that evolves historical atomic positions into their future configurations. Specifically, EGNO (Xu et al., 2024) integrates EGNN and FNO layers to learn dynamic trajectories, capturing spatial and temporal correlations. Recently, transformer neural operators (Bryutkin et al., 2024; Hao et al., 2023; Li et al., 2023a) have surpassed the performance of FNO in most partial differential equation (PDE) tasks. Notably, OFormer (Li et al., 2023a) uses a linear Galerkin-type attention mechanism, which omits the softmax and instead interprets the latent column vectors as basis functions. General Neural Operator Transformer (GNOT) (Hao et al., 2023) employs a novel subquadratic cross-attention methodology to integrate multiple feature types (e.g., shape and point relationships) into their transformer blocks. With ATOM, we unify the MD problem formulation and temporal discretization approach introduced by EGNO with the increased representational power of transformers in operator settings.
@@
MD Benchmarks. Research on graph machine learning for molecular dynamics suffers from poor benchmarking (Bechler-Speicher et al., 2025). For example, despite the fact that MD17 Benzene exhibits non-physical noise approximately 1000 times higher compared to other compounds (Christensen and von Lilienfeld, 2020), it is still regularly employed to benchmark new models (Bihani et al., 2023; Huang et al., 2022; Liao and Smidt, 2023; Xu et al., 2024). The practical relevance of single-task learning on these datasets is also dubious, as predicting trajectories for molecules with existing numerical solutions offers minimal benefit. We believe the strengths of neural approaches emerge in transfer learning, where models generalize to unseen compounds, thereby circumventing the computational costs associated with explicit numerical simulations. This motivates our development of TG80 to facilitate multitask dynamics learning across molecular systems.
@@ 3.1 Problem Formulation
Quasi-equivariance. We formally define quasi-equivariance, motivated by (Elhag et al., 2025).
Definition 3.1 ($\varepsilon$-quasi-equivariance).
We call a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, $\varepsilon$-quasi-equivariant with respect to group $G$ if it satisfies $\mathbb{E}_{x\in\mathcal{X}}\|\int_{G}f(\phi(g)(x))d\mu(g)-\int_{G}\rho(g)(f(x))d\mu(g)\|\leq\varepsilon$, where $\mu$ denotes the normalized Haar measure.
In practice, we approximate the group integration with Monte Carlo samples from $G$.
@@ 3.2 ATOM Model and Training Design
Here we outline the pipeline of ATOM. At its core is an equivariant lifting layer (Section 3.2.1), which maps atomic positions, velocities and their phase features into a richer embedding space while preserving symmetry under the Euclidean group $E(3)$. The lifted embeddings are then processed by the ATOM attention block, which applies heterogeneous attention over positions, velocities, and phase features with chemical augmentation (Section 3.2.2). To capture temporal dynamics, we incorporate a temporal rotary position embedding (T-RoPE) (Section 3.2.2) that depends only on time lags and is shared across atoms, ensuring translation invariance in time and permutation invariance within each molecule. The parameterized ATOM can be written by
$F_{\theta}\coloneqq\mathcal{P}\circ\sigma(\mathcal{K}_{L})\circ\cdots\circ\sigma(\mathcal{K}_{1})\circ\mathcal{Q}$
where $\mathcal{Q}$, $\mathcal{P}$ denotes the equivariant lifting and projection operators respectively. $\mathcal{K}_{l},l=1,...,L$ are the data-dependent kernels induced by cross attention (See Appendix G.1), and $\sigma$ denotes some nonlinear activation function.
Finally, to counter numerical noise in training trajectories, we inject randomly sampled position and velocity perturbations during training (Section 3.2.3), which improves robustness and acts as a regularizer against overfitting. The overall pipeline of ATOM is in Figure 1.
@@ 3.2.1 E(3) Equivariant Lifting
To model atomic states in a symmetry-respecting way, each atom is encoded with its 3D position and velocity, augmented with their norms: $\mathbf{x}=(x,y,z,\sqrt{x^{2}+y^{2}+z^{2}}),\,\mathbf{v}=(v_{x},v_{y},v_{z},\sqrt{v_{x}^{2}+v_{y}^{2}+v_{z}^{2}})$. To construct higher-dimensional features that remain consistent with $E(3)$ symmetry, we apply equivariant lifting that maps the inputs through learnable functions that preserve group actions. Specifically, we use $E(3)$-equivariant linear layers (Geiger and Smidt, 2022) that lift the position and velocity vectors to a feature space wherein they satisfy equivariance constraints by construction. We further construct phase space featrues for each atom by augmenting the position and velocity vectors with atomic number, which is subsequently processed by a learnable equivariant layer to obtain a lifted representation. The final lifted embedding for a molecule is given by $({\mathbf{X}},{\mathbf{V}},{\mathbf{Z}})\in\mathbb{R}^{3\times NP\times d_{v}}$ corresponding to position, velocity and phase features. The second dimension aggregates nodes and time for attention and $d_{v}$ is the embedding space dimension.
We highlight that after the equivariant lifting layer, we do not enforce equivariance in the subsequent transformer blocks. This relaxation improves performance over fully equivariant designs while still showing robustness to random trajectory rotations versus non-equivariant baselines (see Section 4.4).
@@ 3.2.2 ATOM Heterogeneous Temporal Attention
We employ a heterogeneous temporal attention mechanism to enable mixing between multiple features $({\mathbf{X}},{\mathbf{V}},{\mathbf{Z}})\in\mathbb{R}^{3\times NP\times d_{v}}$ across spatial and temporal dimensions. We use the phase space embedding $\mathbf{Z}$} as the query and attend to the key-value pairs formed from all features ${\mathbf{X}},{\mathbf{V}},{\mathbf{Z}}\in\mathbb{R}^{NP\times d_{v}}$. In Figure 6, we show that this improves performance by 6.36% over standard self-attention for single-task prediction. To encode temporal information, we introduce Temporal RoPE (T-RoPE), adapting Rotary Position Embedding (RoPE) (Su et al., 2023) to irregular time lags by scaling the phase rotations according to cumulative timestamps constructed from per-step increments $\Delta t$}.
Let the hidden dimension per head be $d_{h}$ (even). We define frequencies $\omega_{k}=b^{-2k/d_{h}}$ for $k=0,\dots,d_{h}/2-1$. Given per-step time increments $\{\Delta t_{p}\}_{p=1}^{P}$, we build timestamps $t_{p}=t+\sum_{r=1}^{p}\Delta t_{r}$, and assign a single rotation to all $N$ atoms at timestep $p$: $\mathbf{R}_{p}=\operatorname{diag}\left(\mathbf{R}(\theta_{p,0}),\dots,\mathbf{R}(\theta_{p,d_{h}/2-1})\right)\in\mathbb{R}^{d_{h}\times d_{h}}$, where $\theta_{p,k}=\frac{\omega_{k}}{\tau}\,\bigl(t_{p}-t_{0}\bigr)$ and $\mathbf{R}(\theta)\in\mathbb{R}^{2\times 2}$ is the rotation matrix with angle $\theta$ and $\tau>0$ is a timescale hyperparameter. Suppose the query molecule state at time $p$ is given as $\mathbf{Q}_{p}\in\mathbb{R}^{N\times d_{h}}$ and key molecule state at time $p^{\prime}$ is $\mathbf{K}_{p^{\prime}}\in\mathbb{R}^{N\times d_{h}}$. We apply $\mathbf{R}_{p},\mathbf{R}_{p^{\prime}}$ to $\mathbf{Q}_{p},\mathbf{K}_{p^{\prime}}$ respectively so that the rotary dot product $\mathbf{Q}_{p}\mathbf{R}_{p}(\mathbf{K}_{p^{\prime}}\mathbf{R}_{p^{\prime}})^{\top}$ depends only on the time interval $t_{p^{\prime}}-t_{p}$. This makes attention translation invariant in time, which allows for interpolation and extrapolation across irregular increments $\{\Delta t_{p}\}$. In addition, sharing the same $\mathbf{R}_{p}$ across all $N$ atoms in a molecule ensures permutation-invariance within a timestep. For aggregated query and key matrices $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{NP\times d_{h}}$, we denote the application of temporal RoPE across $P$ timesteps and $N$ atoms as $\operatorname{T-RoPE}(\mathbf{Q}),\operatorname{T-RoPE}(\mathbf{K})\in\mathbb{R}^{NP\times d_{h}}$.
@@
where $Q(\cdot),K(\cdot),V(\cdot)$ represent the query, key and value projections. We introduce learnable weights $\gamma_{\mathbf{F}}$ to modulate the relative importance of each feature. In Section G.1, we show that heterogeneous attention is equivalent to a kernel integral operator and discuss its properties.